OpenStudy (anonymous):

Factor the polynomials completely using any method. 4x^3+16x^2-9x-36

OpenStudy (anonymous):

(x+4)(2x-3)(2x+3)

OpenStudy (anonymous):

\[4x^3+16x^2-9x-36\] =\[4x^2(x+4)-9(x+4)\] =\[(4x^2-9)(x+4)\] =\[(2x+3)(2x-3)(x+4)\]